# SpotFormer: Multi-Scale Spatio-Temporal Transformer for Facial Expression Spotting

### v1

@@ Abstract
Facial expression spotting, identifying periods where facial expressions occur in a video, is a significant yet challenging task in facial expression analysis. The issues of irrelevant facial movements and the challenge of detecting subtle motions in micro-expressions remain unresolved, hindering accurate expression spotting. In this paper, we propose an efficient framework for facial expression spotting. First, we propose a Sliding Window-based Multi-Resolution Optical flow (SW-MRO) feature, which calculates multi-resolution optical flow of the input image sequence within compact sliding windows. The window length is tailored to perceive complete micro-expressions and distinguish between general macro- and micro-expressions. SW-MRO can effectively reveal subtle motions while avoiding severe head movement problems. Second, we propose SpotFormer, a multi-scale spatio-temporal Transformer that simultaneously encodes spatio-temporal relationships of the SW-MRO features for accurate frame-level probability estimation. In SpotFormer, our proposed Facial Local Graph Pooling (FLGP) and convolutional layers are applied for multi-scale spatio-temporal feature extraction. We show the validity of the architecture of SpotFormer by comparing it with several model variants. Third, we introduce supervised contrastive learning into SpotFormer to enhance the discriminability between different types of expressions. Extensive experiments on SAMM-LV and CAS(ME)2 show that our method outperforms state-of-the-art models, particularly in micro-expression spotting.
@@ 1 Introduction
Facial expressions can be broadly categorized into two groups: macro-expressions (MaEs) and micro-expressions (MEs). MaEs typically last from 0.5 to 4.0 seconds [1] and are easily perceived by people due to their occurrence on a large facial area and high intensity [2]. The analysis of MaEs is significant in various practical applications, such as sociable robots [3], mental health [4], and virtual reality [5]. In contrast, MEs generally last for less than 0.5 seconds [6], and their perception is much more challenging due to their localized occurrence and low intensity [7]. Because of their involuntary nature, MEs are crucial in situations where people may attempt to conceal emotions or deceive others, such as in lie detection [8], medical care [9], and national security [10]. Therefore, both MaE and ME analysis play important roles in understanding human emotions and behaviors.
@@
In recent years, researchers have made significant progress in developing efficient algorithms for FES. Early studies employed traditional methods to extract hand-crafted features, such as optical flow in large sliding windows. Subsequently, they analyzed feature variations and detected expression intervals using threshold strategies [11, 12, 13, 14]. More recently, with the development of deep learning, more and more researchers introduced learning-based methods to address the problem [15, 16, 17, 18]. Additionally, due to the similarity between FES and temporal action localization (TAL), some researchers applied classical TAL frameworks to FES[19, 20] and introduced additional modules. These methods have achieved considerable improvements in MaE spotting performance.
@@
Despite significant progress made by previous studies, there are still some open questions that require exploration for further performance improvement, especially for ME spotting. The first question is how to extract more robust motion features. Traditional methods typically employ a large sliding window strategy to extract optical flow features and spot potential expression proposals within each sliding window [12, 21]. However, the accuracy of using such a strategy is significantly impacted by head movement issues. In contrast, recent deep learning-based methods calculate optical flow between adjacent frames as motion features [22, 23]. Unfortunately, such an optical flow extraction method fails to effectively reveal subtle motions present in MEs. Therefore, there is an urgent need to develop a strategy that can magnify motion information while mitigating the influence of head movement. Second, some methods [20, 19] applied general TAL frameworks to FES. However, these methods overlook task-specific characteristics, such as the extremely short length of MEs and the problem of irrelevant facial movements. Additionally, they usually employ a pre-trained I3D[24] network to extract features from snippets based on the optical flow extracted between adjacent frames, which is unsuitable for revealing subtle motions as we mentioned. Consequently, their performance, particularly in ME spotting, requires improvement. Third, some methods compute optical flow in specific regions of interest (ROIs) where facial expressions frequently happen to alleviate the influence of irrelevant motions. They then design a network to learn the extracted motion features [23, 18]. However, their proposed networks need a comprehensive consideration of spatio-temporal relationships and multi-scale feature learning. This limitation restricts the representational capacity of their models.
@@
To address the issues mentioned above, we propose an efficient framework for facial expression spotting in untrimmed videos. First, we propose a sliding window-based multi-resolution optical flow (SW-MRO) feature to amplify motion information while mitigating significant head movement problems. The temporal sliding window size is tailored to perceive complete subtle MEs and discern the differences between general MEs and MaEs, enabling our framework to achieve accurate frame-level apex or boundary (onset or offset) probability estimation and expression type classification. Second, we propose SpotFormer, a multi-scale spatio-temporal transformer to capture multi-scale spatial relationships and temporal variations among different facial parts across frames from the SW-MRO features. In SpotFormer, a facial local graph pooling (FLGP) operation is designed to extract multi-scale facial spatial features and learning-based temporal downsampling is employed to extract multi-scale temporal features, enhancing the model’s understanding ability from low-level to high-level. The exploration of several model variants and extensive comparative experiments validate the effectiveness of SpotFormer. Third, we introduce supervised contrastive learning into our model to learn a finer discriminative feature representation for more effective differentiation of various expression types in long videos. This contrastive learning approach mitigates the difficulty in recognizing the boundary to distinguish between specific MaEs.
@@
This paper is an extension of our prior conference publication [25]. The most crucial update is the proposal of SpotFormer. As other differences from the conference paper version, we make a more extensive overview and comparison of the related literature, particularly with regard to TAL. In addition, we present a more comprehensive qualitative and quantitative comparison to previous work in the experimental section. The contributions are listed as follows, where * represents new contributions of this paper:
@@
  • We propose SW-MRO, a sliding window-based multi-resolution optical flow feature for facial expression spotting, which can magnify the motion information while avoiding severe head movement problems. Extensive ablation studies demonstrate that such an optical flow feature can assist the model not only in perceiving complete subtle MEs but also in discerning the differences between general MaEs and MEs.
@@
  • * We present SpotFormer, a novel Transformer-based framework to capture spatial relationships and temporal variations at multiple scales for accurate frame-level probability estimation. In SpotFormer, a facial local graph pooling operation suitable for facial graph structure is proposed for multi-scale spatial feature learning. Extensive comparative experiments on various model architectures and multi-scale feature fusion methods verify the effectiveness of SpotFormer.
@@
  • We introduce supervised contrastive loss to our model for discriminative feature representation learning. To the best of our knowledge, our work is the first to study contrastive learning for facial expression spotting, achieving better recognition of boundaries between different types of expressions.
@@ 2.1 Facial Expression Spotting
In general, FES methods fall into two categories: traditional methods and deep learning methods. Early methods extracted appearance-based features, such as local binary patterns [26] and histogram of oriented gradients [27]. These methods then employed machine learning algorithms for feature difference analysis and utilized threshold strategies for expression spotting. Subsequently, the mainstream approach for motion feature extraction shifted to optical flow. He et al. [12] employed main directional maximal difference analysis [28] to detect facial movements and spot potential expression proposals based on the maximal difference in magnitude along the main direction of optical flow features. He [13] alleviated head movement problems by repeating face alignment using optical flow in the nose tip region. Guo et al. [21] converted optical flow vectors into a polar coordinate representation and introduced a novel decision criterion based on both magnitude and angle information.
@@
With the development of deep learning, an increasing number of researchers have proposed various neural networks for feature learning. Zhang et al. [29] employed convolutional neural networks (CNNs) to extract features from video clips and spotted apex frames by analyzing the feature representations. Verburg et al. [15] extracted histogram of oriented optical flow features to encode the temporal changes in selected facial regions. Then they employed an RNN to spot intervals likely to contain relevant facial movements. Liong et al. [30] introduced a three-stream CNN and employed pseudo-labeling techniques to facilitate the learning process. Yang et al. [31] incorporated facial action unit information and concatenated various types of neural networks for feature learning. Leng et al. [23] extracted several ROIs and adopted the main directional mean optical flow (MDMO) algorithm [32] to compute optical flow between adjacent frames. Then they utilized one-dimensional CNNs to learn temporal variations and estimate the probability of each frame belonging to an apex or boundary frame. Based on [23], Yin et al. [18] learned spatial relations by adding a graph convolutional network (GCN) to embed action unit (AU) label information into the extracted optical flow. Guo et al. [20] proposed a multi-scale local transformer, similar to ActionFormer [33], and considered multi-scale temporal feature fusion for performance improvement. Yu et al. [19] applied A2Net [34] to facial expression spotting and improved it by introducing attention modules. While current optical flow-based methods have significantly improved MaE spotting, their performance in ME spotting remains considerably lower. This is attributed to their extracted optical flow features, which cannot efficiently reveal subtle motions that exist in MEs. To solve this issue, we propose to extract a sliding window-based multi-resolution optical flow feature, which can effectively magnify these subtle motions and enhance the distinction between MaEs and MEs, thereby achieving accurate frame-level probability estimation and improving MaE and ME spotting performance.
@@ 2.2 Temporal Action Localization
Similar to FES, TAL aims to detect action intervals in general scene videos. Deep learning-based methods for TAL can be divided into anchor-based methods and anchor-free methods, with the latter further divided into frame-level probability estimation methods and boundary regression-based methods.
@@
Early works treated TAL as a 1D object detection task. Xu et al. [35] proposed a region convolutional 3D network to encode the video streams and locate the action proposals using an anchor-based method. Chao et al. [36] refined the faster R-CNN object detection framework to make it suitable for TAL, employing a multi-scale architecture to accommodate extreme variations in action duration. However, anchor-based methods have some drawbacks. First, they require pre-defining a fixed number of anchors, resulting in increased computation and memory requirements. Second, the length and number of anchors need careful pre-definition, limiting flexibility and affecting the generalization ability of the proposed model.
In 2018, Lin et al. [37] proposed a Boundary-Sensitive Network (BSN), pioneering a method that first locates temporal boundaries with high probabilities and then aggregates these probabilities to generate action proposals. This is the first frame-level probability estimation method, and it then became the mainstream. Subsequently, Lin et al. [38] refined BSN [37] by proposing a boundary-matching network, incorporating a boundary-matching mechanism to evaluate confidence scores of densely distributed proposals. Zhao et al. [39] argued that frame-level probability-based methods had a limited understanding of the temporal dimension, leading to inconsistency and discontinuity. They addressed this issue by proposing two consistency losses to mutually regularize the learning process. Recently, researchers have focused on refining proposals generated by previous methods. Zhu et al. [40] refined proposals generated by BSN [37] by introducing a ContextLoc, which models the local context, global context, and context-aware inter-proposal relations in a unified framework. Nag et al. [41] presented a novel Gaussian aware post-processing method for more accurate model inference, which models the start and end points of action instances with a Gaussian distribution to enable temporal boundary inference at a sub-snippet level.
@@
Yang et al. [34] were the first to directly regress the boundary offset for each temporal point, which solves the inflexibility and computation problems of anchor-based methods. Zhang et al. presented ActionFormer [33], a simple yet effective multi-scale transformer designed to further improve action localization performance. Shao et al. [42] proposed an action sensitivity learning framework, which aims to assess the value of each frame and leverage the generated action sensitivity to recalibrate the training procedure.
@@
Even though there are many similarities between TAL and FES, there are still some differences that make the direct application of TAL methods into FES inappropriate. The main differences can be summarized in three aspects. Firstly, motion information is usually obvious to detect in TAL videos, while in FES, the main challenge is to capture the subtle motion of MEs and suppress noises such as head movement and eye blinking. Since anchor-based methods and boundary regression-based methods typically require a large receptive field, as discussed before, extracting delicate optical flow becomes challenging. Therefore, in this paper, we design a frame-level probability-based framework.
@@
Secondly, the scene usually changes in TAL videos, and the main challenge is to distinguish between the action scene and background change. In contrast, the scene during the whole video does not change much in FES, making it easier to capture vital motion information in certain facial areas that are strongly related to facial expressions. Therefore, in this paper, we propose a graph-based method to efficiently extract motion features while suppressing noises.
@@
Finally, TAL involves an action classification task for each proposal since their videos may involve hundreds of action classes. However, there are only two action classes (i.e., MaE and ME) in FES. Rather than focusing on action recognition, we aim to distinguish between MEs and MaEs. Therefore, in this paper, we emphasize perceiving MEs and analyzing the differences between MEs and MaEs within compact sliding windows.
@@
Figure 3: Extracted ROIs and constructed facial graph structure are denoted in yellow, while the nose tip region for face alignment is denoted in green. Note that the edge connections are only for presentation, they are not involved in the model.

### v2

@@ Abstract
Facial expression spotting, identifying periods where facial expressions occur in a video, is a significant yet challenging task in facial expression analysis. The issues of irrelevant facial movements and the challenge of detecting subtle motions in micro-expressions remain unresolved, hindering accurate expression spotting. In this paper, we propose an efficient framework for facial expression spotting. First, we propose a Sliding Window-based multi-temporal-resolution Optical flow (SW-MRO) feature, which calculates multi-temporal-resolution optical flow of the input image sequence within compact sliding windows. The window length is tailored to perceive complete micro-expressions and distinguish between general macro- and micro-expressions. SW-MRO can effectively reveal subtle motions while avoiding the optical flow being dominated by head movements. Second, we propose SpotFormer, a multi-scale spatio-temporal Transformer that simultaneously encodes spatio-temporal relationships of the SW-MRO features for accurate frame-level probability estimation. In SpotFormer, we use the proposed Facial Local Graph Pooling (FLGP) operation and convolutional layers to extract multi-scale spatio-temporal features. We show the validity of the architecture of SpotFormer by comparing it with several model variants. Third, we introduce supervised contrastive learning into SpotFormer to enhance the discriminability between different types of expressions. Extensive experiments on SAMM-LV, CAS(ME)2, and CAS(ME)3 show that our method outperforms state-of-the-art models, particularly in micro-expression spotting. Code will be available at https://github.com/KinopioIsAllIn/SpotFormer.
@@ 1 Introduction
Facial expressions can be broadly categorized into two groups: macro-expressions (MaEs) and micro-expressions (MEs). MaEs typically last from 0.5 to 4.0 seconds [nummenmaaekman] and are easily perceived by people due to their occurrence on a large facial area and high intensity [corneanu2016survey]. The analysis of MaEs is significant in various practical applications, such as sociable robots [fukuda2002facial], mental health [kopper1996experience], and virtual reality [facialvr]. In contrast, MEs generally last for less than 0.5 seconds [yan2013fast], and their perception is much more challenging due to their localized occurrence [bhushan2015study] and low intensity [wang2025micro]. Because of their involuntary nature, MEs are crucial in situations where people may attempt to conceal emotions or deceive others, such as in lie detection [ekman2009telling], medical care [endres2009micro], and national security [o2009police]. Therefore, both MaE and ME analysis play important roles in understanding human emotions and behaviors.
@@
In recent years, researchers have made significant progress in developing efficient algorithms for FES. Early studies employed traditional methods to extract hand-crafted features, such as optical flow in large sliding windows. Subsequently, they analyzed feature variations and detected expression intervals using threshold strategies [gan2020optical, he2020spotting, yuhong2021research, zhao2022rethinking]. More recently, with the development of deep learning, more and more researchers introduced learning-based methods to address the problem [verburg2019micro, wang2021mesnet, yap20223d, yin2023aware]. Additionally, due to the similarity between FES and temporal action localization (TAL), some researchers applied classical TAL frameworks to FES[yu2023lgsnet, GUO2023146] and introduced additional modules. These methods have achieved considerable improvements in MaE spotting performance.
@@
Despite significant progress made by previous studies, there are still some open questions that require exploration for further performance improvement, especially for ME spotting. The first question is how to extract more robust motion features. Traditional methods typically employ a large sliding window strategy to extract optical flow features and spot potential expression proposals within each sliding window [he2020spotting, guo2021magnitude]. However, the accuracy of using such a strategy is significantly impacted by head movement issues. In contrast, recent deep learning-based methods calculate optical flow between adjacent frames as motion features [yu2021lssnet, leng2022abpn]. Nevertheless, such an optical flow extraction method is insufficient to effectively reveal subtle motions present in MEs. Therefore, there is an urgent need to develop a strategy that can magnify motion information while mitigating the influence of head movement. Second, some methods [GUO2023146, yu2023lgsnet] applied general TAL frameworks to FES and refined them by introduing novel modules. However, these methods usually employ a pre-trained I3D[carreira2017quo] network to extract features from snippets based on the optical flow extracted between adjacent frames, which is unsuitable for revealing subtle motions as we mentioned before. Consequently, their performance, particularly in ME spotting, requires improvement. Third, some methods compute optical flow in specific regions of interest (ROIs) where facial expressions frequently happen to alleviate the influence of irrelevant motions. They then design a network to learn the extracted motion features [leng2022abpn, yin2023aware]. However, their proposed networks need a more comprehensive consideration of spatio-temporal relationships and multi-scale feature learning. This limitation restricts the representational capacity of the models.
@@
To address the issues mentioned above, we propose an efficient framework for facial expression spotting in untrimmed videos. First, we propose a sliding window-based multi-temporal-resolution optical flow (SW-MRO) feature to amplify motion information while mitigating head movement problems. Specifically, we calculate the optical flow between the first frame and each subsequent frame within compact sliding windows to generate multi-temporal-resolution optical flow features. Moreover, the temporal sliding window size is tailored to perceive complete subtle MEs and discern the differences between general MEs and MaEs. The SW-MRO feature enables our framework to achieve accurate frame-level apex or boundary (onset or offset) probability estimation and expression type classification. Second, we propose SpotFormer, a multi-scale spatio-temporal transformer to capture multi-scale spatial relationships and temporal variations among different facial parts across frames from the SW-MRO features. In SpotFormer, a facial local graph pooling (FLGP) operation is designed to extract multi-scale facial spatial features, while a learning-based temporal downsampling strategy is employed to extract multi-scale temporal features. These designs enhance the model’s understanding ability from low-level to high-level. The exploration of several model variants and extensive comparative experiments validate the effectiveness of SpotFormer. Third, we introduce supervised contrastive learning into our model to learn a finer discriminative feature representation. This contrastive learning approach alleviates the challenge of recognizing the classification boundaries between MaEs, MEs, and neutral expressions at the frame level.
@@
This paper is an extension of our prior conference publication [deng2024SpoTGCN]. The most crucial update is the proposal of SpotFormer. As other differences from the conference paper version, we make a more extensive overview and comparison of the related literature, particularly with regard to TAL. In addition, we present a more comprehensive qualitative and quantitative comparison to previous work in the experimental section. The contributions are listed as follows, where * represents new contributions of this paper:
@@
  • We propose SW-MRO, a sliding window-based multi-temporal-resolution optical flow feature for facial expression spotting, which can magnify the motion information while avoiding the optical flow being dominated by head movements. Extensive ablation studies demonstrate that such a carefully designed optical flow feature can assist the model not only in perceiving complete subtle MEs but also in discerning the differences between general MaEs and MEs.
@@
  • * We present SpotFormer, a novel Transformer-based framework to capture spatial relationships and temporal variations at multiple scales for accurate frame-level probability estimation. In SpotFormer, we propose a facial local graph pooling operation tailored to the facial graph structure for multi-scale spatial feature learning, while learnable convolution layers are employed for multi-scale temporal feature learning. Extensive comparative experiments on various model architectures verify the effectiveness of SpotFormer.
@@
  • We introduce supervised contrastive learning into our model to enhance the discriminative feature representation. To the best of our knowledge, our work is the first to study contrastive learning for facial expression spotting, which enables our model to learn better classification boundaries between different expression categories at the frame level.
@@ 2.1 Facial Expression Spotting
Early FES methods extracted appearance-based features, such as local binary patterns [moilanen2014spotting] and histogram of oriented gradients [davison2015micro]. These methods then employed machine learning algorithms and threshold strategies for expression spotting. Subsequently, See et al. [see2019megc] proposed the Facial Micro-Expressions Grand Challenge (MEGC), which introduced unified evaluation datasets and metrics for the FES task. The subsequent MEGC series [jingting2020megc2020, li2022megc2022, davison2023megc2023, see2024megc2024] have attracted increasing attention from researchers and significantly promoted the development of this field.
@@
In recent years, the mainstream approach for motion feature extraction shifted to optical flow. He et al. [he2020spotting] employed main directional maximal difference analysis [wang2017main] to detect facial movements and spot potential expression proposals. He [yuhong2021research] alleviated head movement problems by repeating face alignment using optical flow in the nose tip region. Guo et al. [guo2021magnitude] converted optical flow vectors into a polar coordinate representation and introduced a novel decision criterion accordingly. Recently, several researchers [zhao2022rethinking, yu2024micro] proposed novel methods to refine optical flow features or expression proposals, which further improved performance.
@@
With the development of deep learning, an increasing number of researchers have proposed various neural networks for feature learning. Zhang et al. [zhang2018smeconvnet] employed convolutional neural networks (CNNs) to extract features from video clips and spotted apex frames by analyzing the feature representations. Verburg et al. [verburg2019micro] extracted histogram of oriented optical flow features and employed an RNN to spot expression intervals. Liong et al. [liong2021shallow] introduced a three-stream CNN and employed pseudo-labeling techniques to facilitate the learning process. Yang et al. [yang2021facial] incorporated facial action unit information and concatenated various types of neural networks for feature learning. Leng et al. [leng2022abpn] calculated the main directional mean optical flow (MDMO) [liu2015main] optical flow features between adjacent frames and extended BSN [Lin_2018_ECCV] for frame-level probability estimation. Yin et al. [yin2023aware] extended ABPN [leng2022abpn] by introducing a graph convolutional network (GCN) to embed action unit (AU) label information into the extracted optical flow. Guo et al. [GUO2023146] proposed a multi-scale local transformer, similar to ActionFormer [zhang2022actionformer], and considered multi-scale temporal feature fusion for performance improvement. Yu et al. [yu2023lgsnet] applied A2Net [yang2020revisiting] to facial expression spotting and improved it by introducing attention modules. Zhang et al. [zhang2024multi] proposed a framework that integrates VideoMAE-based feature extraction, multi-scale candidate segment generation, and a multi-start-point optical flow filtering method for accurate ME and MaE detection.
While current optical flow-based methods have significantly improved MaE spotting, their performance in ME spotting remains considerably lower. This is attributed to their extracted optical flow features, which cannot efficiently reveal subtle motions that exist in MEs. To solve this issue, we propose to extract a sliding window-based multi-temporal-resolution optical flow feature, which can effectively magnify these subtle motions and enhance the distinction between MaEs and MEs, thereby achieving accurate frame-level probability estimation and improving MaE and ME spotting performance.
@@ 2.2 Temporal Action Localization
Similar to FES, TAL aims to detect action intervals in general scene videos. Early works treated TAL as a 1D object detection task. Xu et al. [Xu_2017_ICCV] proposed a region convolutional 3D network to encode the video streams and locate the action proposals using an anchor-based method. Chao et al. [Chao_2018_CVPR] applied the faster R-CNN object detection framework to TAL by employing a multi-scale architecture to accommodate extreme variations in action duration. However, anchor-based methods require pre-defining a fixed number of anchors. The length and number of anchors need careful pre-definition, limiting flexibility and affecting the generalization ability of the proposed model.
In 2018, Lin et al. [Lin_2018_ECCV] proposed the Boundary-Sensitive Network (BSN), which was the first method to estimate frame-level probabilities and subsequently became the mainstream approach. Recently, Nag et al. [Nag_2023_CVPR] presented a novel Gaussian aware post-processing method for more accurate model inference, which models the start and end points of action instances with a Gaussian distribution to enable temporal boundary inference at a sub-snippet level.
@@
Yang et al. [yang2020revisiting] were the first to directly regress the boundary offset for each temporal point, which solves the inflexibility and computation problems of anchor-based methods. Zhang et al. presented ActionFormer [zhang2022actionformer], a simple yet effective multi-scale transformer designed to further improve action localization performance.
Even though there are many similarities between TAL and FES, there are still some differences between the two tasks that may limit the performance improvements when directly applying TAL methods to FES.
Here, we summarize three main differences that motivate us to consider the specific characteristics of the FES task more thoroughly when developing our unique method.
@@
Firstly, motion information is usually obvious to detect in TAL videos, while in FES, the main challenge is to capture the subtle motion of MEs and suppress noises such as head movement and eye blinking. Since anchor-based methods and boundary regression-based methods typically require a large receptive field, as discussed before, extracting delicate optical flow becomes challenging. Therefore, in this paper, we design a frame-level probability-based framework following BSN [Lin_2018_ECCV].
@@
Secondly, the scene usually changes in TAL videos, and the main challenge is to distinguish between the action scene and background change. In contrast, the scene during the whole video does not change much in FES, making it easier to capture vital motion information in certain facial areas that are strongly related to facial expressions. Therefore, in this paper, we represent the SW-MRO feature as spatio-temporal facial graphs and propose the SpotFormer to efficiently analyze the latent motion patterns.
@@
Finally, TAL involves an action classification task for each proposal since their videos may involve hundreds of action classes. However, there are only two action classes (i.e., MaE and ME) in FES. Rather than focusing on action recognition, we aim to distinguish between MEs and MaEs. Therefore, in this paper, we emphasize perceiving MEs and analyzing the differences between MEs and MaEs within compact sliding windows, which inspired us to propose the SW-MRO feature and the supervised contrastive learning.
@@
Figure 3: Extracted ROIs and constructed facial graph structure are denoted in yellow, while the nose tip region for face alignment is denoted in green.